# Search for Long-lived Particles at Future Lepton Colliders Using Deep Learning Techniques

### v2

@@ IV Acknowledgement
The authors thank theorists Zhen Liu, Tao Liu, Jia Liu, Ziping Wang, as well as experimentalists Manqi Ruan and Gang Li for their useful suggestions and inputs about the paper. This work was sponsored by the National Key R&D Program of China (2022ZD0117804), National Natural Science Foundation of China (Grants No. 12211540001 and No. 12305217), Key Research Project of Zhejiang Lab (No. 2022PI0AC01), Guangdong Major Project of Basic and Applied Basic Research (No. 2020B0301030008), Science and Technology Program of Guangzhou (No. 2019050001).
Supplementary Information
I.1 A selection-based analysis
@@
Searching for LLPs with the selection-based method is roughly carried out using various reconstruction observable. However, we only investigate a special case that LLPs have long displaced vertices and decay inside the muon detector. In this case, lots of shower hits and an energy burst are expected in the muon detector. Thus, we can save most signal events while almost kicking out all SM background events after applying the following selections,
@@
  • $\Delta T_{j}=\min\left(t_{\text{hit},i}-r_{\text{hit},i}/c\right)\geq 3$ ns: the minimal time difference, where $t_{\text{hit},i}$ represents the hitting time of the $i^{\text{th}}$ component in the jet cluster measured by the muon spectrometer and $r_{\text{hit},i}$ is the $i^{\text{th}}$ Euclidean distance to IP, and c the light speed in vacuum.
@@
Traditional methods generally faces various difficulties in reconstructing and identifying both Z bosons and LLPs, and thus require complicated case-by-case analyses, while the machine learning approach does not. The latter approach uniformly makes use of all available information of event data, and not only exhibits a technical superiority but also systematically improves signal efficiencies, as shown in Table 1 in the paper and Table 5. The signal efficiency is about 97%-98% with a LLP lifetime of 10 ns and mass of 50 GeV, significantly higher than 50.8% obtained with the traditional approach.
Table 5: Number of signal and background events after various selection criteria. MC simulation samples are produced corresponding to the integrated luminosity 5.6 $ab^{-1}$. Signal events have a LLP lifetime of 20 ns and mass of 50 GeV.
@@
I.2 Searching for LLPs by placing an external detector
We explore the potential of enhancing the discovery sensitivity by placing an external detector far away from the baseline detector. A typical external detector is a stack consisting of multi-layer scintillators. The distance between the outermost and innermost layers is 100 meters. This detector only covers the barrel region, corresponding to an approximated geometry acceptance of 0.65, which allows us to detect LLPs decaying outside the baseline detector from 6 meters to 106 meters. The gain factor is roughly evaluated in accordance with signal efficiency linked to geometry parameters of the external detector,
$F_{\textrm{gain}}=\frac{N_{\textrm{obs}}}{N_{\textrm{gen}}}=\frac{\Delta\Omega}{4\pi}\frac{\Delta L}{d}e^{-\frac{L}{d}}$ (2)
where $N_{obs}$ represents the number of observed events among the number of generated events expected, $\Delta\Omega$ represents overall acceptance of external detector among $4\pi$ solid space, $\Delta L$ the longitudinal length of external detector, and d the expected decay length of LLPs in Lab. frame. Table 6 summarizes gain factors compared to the baseline detector. A maximum gain factor of 16.2 can be achieved.
Table 6: The expected gain factor by placing an external detector in detecting LLPs with various masses and lifetimes.
@@ I.2 Searching for LLPs by placing an external detector
Ext. Detector 1 1 1 3.2 11.6 16.2
10 1 1 1 3.3 11.8
50 1 1 1 1.1 3.6
I.3 Training Loss
I.4 Two dimensional upper limits on ($\mathcal{B}_{\textrm{2-jet}}$, $\mathcal{B}_{\textrm{4-jet}}$)

### v5

@@ 7 Conclusion
Acknowledgement The authors thank theorists Zhen Liu, Tao Liu, Jia Liu, Ziping Wang, as well as experimentalists Manqi Ruan and Gang Li for their useful suggestions and inputs about the paper. This work was supported by National Key R&D Program of China (Grant No. 2024YFA1610603, 2023YFA1606003); National Natural Science Foundation of China (Grants No. 12305217, 12375072 and 12475108).
Appendix B Pile-up and cosmic ray backgrounds
Unlike hadron colliders, lepton colliders typically exhibit a significantly lower level of pile-up due to the cleaner nature of lepton interactions and the generally lower total cross-sections involved. The pile-up ($\mu$) at a lepton collider can be quantitatively estimated using the formula $\mu=L\cdot\sigma\cdot\tau$, where $L$ is the luminosity, $\sigma$ is the total cross-section, and $\tau$ is the bunch spacing.
Given the parameters for the CEPC [42], with $L=5\times 10^{34}\,\textrm{cm}^{-2}\textrm{s}^{-1}$, $\sigma=1000\times 10^{-36}\,\textrm{cm}^{2}$, and $\tau=636\times 10^{-9}\,\textrm{s}$, the computed pile-up rate is approximately $3\times 10^{-5}$. This low pile-up rate underscores the minimal impact of pile-up on the CEPC’s operation, making the pile-up background negligible.
The cosmic ray background, particularly from muons, is effectively mitigated in lepton colliders through several ways, making it a negligible concern in most analyses. Firstly, the timing method uses the fact that cosmic muons do not coincide with the precise timing of collision events, allowing for a clear differentiation based on the synchronization of particle detection with beam crossings. Secondly, the distinct trajectories of cosmic muons, which are typically vertical, are different from the varied directions of muons originating from collision points. Moreover, the underground placement of detectors significantly shields them from cosmic interference, complemented by additional physical barriers that absorb cosmic radiation. The combination of these strategies effectively minimizes the cosmic ray background noise. This background is also ignored in other LLPs search experiments such as the MATHUSLA experiment [56].
Appendix C Training losses
Appendix D Sensitivity study on $\mathcal{B}(H\rightarrow XX)$
Figures 10 and 11 present the confidence limits derived from pseudo-experiments generated under the null hypothesis, as functions of the signal strength parameter $\mu$. The parameter $\epsilon_{V}$, defined as the ratio of branching fractions $\epsilon_{V}=\frac{BR(X\to\nu\bar{\nu})}{BR(X\to q\bar{q})}$, is either fixed to 0.2 (Figure 10) or allowed to float freely during the fitting process (Figure 11).
Figure 10: Confidence limits obtained from pseudo-experiments under the null hypothesis, as a function of the signal strength ($\mu$), with a fixed ratio $\epsilon_{V}=\frac{BR(X\to\nu\bar{\nu})}{BR(X\to q\bar{q})}=0.2$. Columns correspond to LLP masses of 50, 10, and 1 GeV, respectively (left to right), and rows correspond to LLP lifetimes of 0.0001, 0.1, 1.0, 10, and 100 nanoseconds (top to bottom).
In both figures, the LLP mass scenarios vary horizontally (from left to right: 50, 10, and 1 GeV), while LLP lifetimes vary vertically (from top to bottom: 0.0001, 0.1, 1.0, 10, and 100 nanoseconds).
Figure 11: Confidence limits obtained from pseudo-experiments under the null hypothesis, as a function of the signal strength ($\mu$), with the ratio $\epsilon_{V}=\frac{BR(X\to\nu\bar{\nu})}{BR(X\to q\bar{q})}$ allowed to float during fitting. Columns correspond to LLP masses of 50, 10, and 1 GeV, respectively (left to right), and rows correspond to LLP lifetimes of 0.0001, 0.1, 1.0, 10, and 100 nanoseconds (top to bottom).
Appendix E Two dimensional upper limits on ($\mathcal{B}_{\textrm{2-jet}}$, $\mathcal{B}_{\textrm{4-jet}}$)
Appendix F Sensitivity with different background estimation
Both 1-D and 2-D exclusion limits are summarized in Table 6, but with 2 backgrounds.
Table 6: The 95% C.L. exclusion limit on BR($h\to XX$) for all signal channels with both fixed and floating $\epsilon_{V}$ with 2 backgrounds assumption. The limits include ±$\pm 1\sigma$ uncertainties after taking into account both statistical and systematic contributions.
Scenario $\mathcal{B}$ (×$\times 10^{-6}$) Lifetime [ns]
Mass [GeV] 0.001 0.1 1 10 100
Fixed 1 $2.78^{+1.21}_{-0.54}$ $2.40^{+0.95}_{-0.55}$ $5.21^{+1.96}_{-1.47}$ $34.60^{+12.43}_{-10.17}$ $1285.89^{+496.57}_{-301.57}$
10 $3.58^{+1.44}_{-0.87}$ $1.74^{+0.75}_{-0.42}$ $2.07^{+0.79}_{-0.50}$ $4.42^{+1.68}_{-1.06}$ $31.94^{+11.97}_{-8.64}$
50 $1.84^{+0.74}_{-0.44}$ $1.54^{+0.59}_{-0.35}$ $1.68^{+0.64}_{-0.40}$ $1.69^{+0.64}_{-0.40}$ $4.04^{+1.54}_{-1.02}$
Floating 1 $5.62^{+15.39}_{-4.07}$ $2.63^{+4.82}_{-1.40}$ $2.80^{+11.86}_{-0.87}$ $17.54^{+82.75}_{-5.21}$ $675.25^{+3739.15}_{-195.72}$
10 $3.75^{+11.82}_{-1.86}$ $4.04^{+10.78}_{-3.08}$ $1.08^{+4.85}_{-0.31}$ $2.30^{+10.34}_{-0.65}$ $16.74^{+76.53}_{-4.81}$
50 $3.58^{+11.01}_{-2.59}$ $2.78^{+8.31}_{-1.93}$ $1.96^{+3.24}_{-1.10}$ $1.97^{+3.27}_{-1.10}$ $2.15^{+9.64}_{-0.71}$
Appendix G Comparison with selection-based analyses
@@
LLPs searches with the selection-based method has also been carried out. For simplicity, we have investigated a special case that LLPs have long displaced vertices and decay inside the muon detector. In this case, shower hits and energy bursts from LLPs decay products are expected in the muon detector while SM backgrounds deposit most energy and hits in the inner detectors. Almost all SM backgrounds are rejected after applying the following selections,
@@
  • $\Delta T_{j}=\min\left(t_{\textrm{hit},i}-r_{\textrm{hit},i}/c\right)\geq 3$ ns: the minimal time difference, where $t_{\textrm{hit},i}$ represents the hitting time of the $i^{\textrm{th}}$ component in the jet cluster measured by the muon spectrometer and $r_{\textrm{hit},i}$ is the $i^{\textrm{th}}$ Euclidean distance to IP, and c the light speed in vacuum.
@@
Traditional methods generally faces various difficulties in reconstructing and identifying both Z bosons and LLPs, and thus require complicated case-by-case analyses, while the machine learning approach does not. The latter approach uniformly makes use of all available information of event data, and not only exhibits a technical superiority but also systematically improves signal efficiencies. Comparing Table 4 and Table 7, the ML-based signal efficiency is about 97%-98% with a LLP lifetime of 10 ns and mass of 50 GeV, significantly higher than 50.8% obtained with the selection-based approach.
We have also compared our results with a previous LLP search [57] conducted at lepton colliders, where the selection-based method is implemented with fine-tuning and optimization for varying scenarios, achieving a maximum selection efficiency ***”Selection efficiency” specifically denotes the product of signal acceptance and signal efficiency. of 0.76 for a 50 GeV particle with a 1 ns lifetime. In contrast, our ML approach significantly outperforms this result and achieves a maximum efficiency of 0.99. Notably, our method maintained high selection efficiencies (over 0.9) for lifetimes under 1 ns and masses above 10 GeV, and reasonable efficiencies (around 0.4) for masses as low as 1 GeV. Furthermore, while the selection-based analysis is limited to lifetimes of up to 1 ns, our ML approach can reach up to 100 ns, demonstrating its superior capability in performing a broader spectrum of LLP searches.
In summary, we have developed an ML-based approach for LLP searches that outperforms traditional selection-based methods on almost all fronts.
Table 7: Number of signal and background events after various selection criteria. MC simulation samples are produced corresponding to the integrated luminosity 5.6 $ab^{-1}$. Signal events have an LLP lifetime of 20 ns and a mass of 50 GeV.
@@
Appendix H Searching for LLPs with an external detector
We investigate the possibility of enhancing the sensitivity for detecting LLPs by incorporating an external detector placed significantly farther from the interaction point than the baseline detector. For example, such an external detector may consist of stacked, multi-layer scintillator arrays in the barrel region outside the baseline detector, named as the Far Barrel Detector (FBD), designed to detect LLP decays occurring outside the primary detector. To quantify the sensitivity gain by placing an external detector, we define a gain factor $F_{\textrm{gain}}$ as follows:
$F_{\textrm{gain}}=\frac{N_{\textrm{obs}}}{N_{\textrm{gen}}}=\frac{\Delta\Omega}{4\pi}\left(\frac{e^{-R_{\textrm{max}}/d}-e^{-(R_{\textrm{max}}+\Delta L)/d}}{1-e^{-R_{\textrm{max}}/d}}\right)+1,$ (2)
where $N_{\textrm{obs}}$ denotes the number of LLP events observed in the FBD, $N_{\textrm{gen}}$ is the total number of LLP events generated, and $\Delta\Omega$ is the geometric acceptance of the external detector relative to the full $4\pi$ solid angle. Here, $R_{\textrm{max}}$ is the outer radius of the baseline detector (beyond which the LLPs enter the FBD), $d$ is the decay length of the LLPs in the laboratory frame, and $\Delta L$ is the gap between the FBD and the outer boundary of the baseline detector.
Figure 13: Layout of the Far Barrel Detector (solid lines) relative to the CEPC baseline detector (red) in the $x$-$y$ (left) and $y$-$z$ (right) planes.
Table 8: Sensitivity gain factor $F_{\textrm{gain}}$ estimated for different LLP masses and lifetimes.
@@ Appendix H Searching for LLPs with an external detector
1 1 1 2.8 9.9 13.7
10 1 1 1 2.9 10.1
50 1 1 1 1.1 3.3
Figure 13 illustrates the relative positions of the external and baseline detectors, where $R_{\textrm{max}}$ is 6 meters and $\Delta L$ is 100 meters. The estimated sensitivity gain factor $F_{\textrm{gain}}$ obtained by adding the external detector in the barrel region with a geometry acceptance factor of about 0.7 is summarized in Table 8, highlighting significant improvements, especially for lighter LLPs and longer lifetimes where a maximum gain factor of 13.7 can be achieved. Such an external detector setup thus offers valuable complementary sensitivity to LLP searches at future collider experiments.